Hydrochloric acid and distilled water were mixed to produce 210mL of solution. If 120mL more H2O was used than HCI, how many mL of each were mixed?
tnx!.. pls answer as soon as possible!..
tnx!

Answer by venugopalramana(3286)

Hydrochloric acid and distilled water were mixed to produce 210mL of solution. If 120mL more H2O was used than HCI, how many mL of each were mixed?
tnx!.. pls answer as soon as possible!..
tnx!
let hcl used =x ml.
120 more =x+120 ml.
total mix =x+x+120=2x+120=210
2x=210-120=90
x=90/2=45 ml.
hcl= 45 ml
water = 45+120=165 ml.